Ammonia NH3 chemically reacts with oxygen gas O2 to produce nitric oxide NO and water H2O.

What mass of water is produced by the reaction of 2.1g of oxygen gas?


1
Expert's answer
2021-03-17T06:37:15-0400

5O2 + 4NH3 = 6H2O + 4NO

5O2 = 6H2O

5 × 32g O2 = 6 × 18g H2O

1g O2 = 6×18/(5×32)g H2O

2.1g O2 = 6×18×2.1/(5×32) g H2O

= 1.42g H2O